OpenStudy (anonymous):

sin3x cosx - cos3x sinx

OpenStudy (solomonzelman):

Is it, \(\Large\color{blue}{ \sin^3x\cos x - \cos^3x\sin x }\) and you want to smplify it?

OpenStudy (anonymous):

sinAcosB-cosAsinB=sin(A-B)